# Learning 3D Representations from Procedural 3D Programs

### v2

@@ 4.5 Analysis
Fig. 4 displays samples from each configuration alongside quantitative comparisons of pretraining performance and downstream classification accuracy on PB-T50-RS, the most challenging variant of ScanObjectNN [34]. As shape complexity increases, the pretraining task becomes more difficult, leading to higher reconstruction losses at the 300th training epoch. However, the downstream classification performance of Point-MAE-Zero improves. This underscores the importance of topological diversity in shapes for effective self-supervised point cloud representation learning.
@@
Figure 5: Impact of pretraining dataset size. We report the classification accuracy (%) on the PB-T50-RS subset of ScanObjectNN [34] as a function of the pretraining dataset size.
@@
Dataset Size. Fig. 5 illustrates the effect of dataset size (i.e., the number of procedurally generated 3D shapes) on Point-MAE-Zero’s performance in the shape classification task on the PB-T50-RS benchmark.
Our experiments show that performance improves as dataset size increases, despite the dataset being procedurally generated. However, simply enlarging the dataset appears to yield diminishing returns, which may be due to intrinsic limitations of a purely synthetic 3D dataset or the representation learning bottleneck within Point-MAE.
Notably, Point-MAE-Zero and Point-MAE-SN achieve comparable performance on downstream tasks when pretrained on datasets of the same size, regardless of differences in their pretraining data domains (i.e., synthetic shapes vs. ShapeNet). We include additional experiments in the appendix to ablate the effect of dataset size.
Efficiency of Transfer Learning. Fig. 6 shows the learning curves for training from scratch, Point-MAE-SN, and Point-MAE-Zero on shape classification tasks in the transfer learning setting. Both Point-MAE-SN and Point-MAE-Zero demonstrate faster training convergence and higher test accuracy compared to training from scratch. This trend is consistent across both ModelNet40 and ScanObjectNN benchmarks.
t-SNE Visualization. Fig. 7 visualizes the distribution of 3D shape representations from Point-MAE-SN and Point-MAE-Zero via t-SNE [35], before and after fine-tuning on specific downstream tasks. It also includes representations from a randomly initialized neural network as a reference.
First, compared to the representations from scratch, both Point-MAE-SN and Point-MAE-Zero demonstrate visually improved separation between different categories in the latent space. For example, this is evident in the red and light blue clusters on ModelNet40 and the blue and light blue clusters on ScanObjectNN. This indicates the effectiveness of self-supervised 3D representation learning via masked auto-encoding.
@@ 5 Discussion
This work has several limitations. For example, due to limited computational resources, we were unable to further scale up our experiments, such as by increasing the dataset size or conducting more detailed ablation studies on procedural 3D generation. Additionally, our findings may be influenced by potential biases in visualization tools (e.g., t-SNE) or benchmarks (e.g., data distribution and evaluation protocols). Furthermore, in 3D vision, the distinction between geometric structures and semantics remains an open question, as well-stated by Xie et al. [45]. This work also does not provide any novel representation learning method. Nevertheless, we hope our findings will inspire further exploration into self-supervised 3D representation learning.
6 Acknowledgement
The authors gratefully acknowledge Research Computing at the University of Virginia for providing the computational resources and technical support that made the results in this work possible. Research Computing at UVA.
@@ 6 Acknowledgement
We provide more implementation details, additional experimental results, and visualizations in this supplementary material. In Sec. A, we present our training-from-scratch baseline, which surpasses the results reported in Point-MAE [26]. In Sec. B, we introduce a more rigorous evaluation protocol using a dedicated validation set instead of relying on the test set for validation. Moreover, in Sec. C, we provide linear probing results on ModelNet40 and three ScanObjectNN variants to compare the performance of Point-MAE-SN and Point-MAE-Zero. Additional visualizations are provided in Sec.D. Details on generating scene-level procedural 3D programs are in Sec.E. Lastly, implementation details for 3D object SSLs and 3D scene SSLs are presented in Sec. F.
@@ Appendix A Training-from-Scratch Baseline
For the training-from-scratch baseline, we report the results in the supplementary material for completeness. In the main text, we referenced the scores reported in the original Point-MAE paper for this baseline. However, our experiments suggest that training from scratch in our setup produced higher scores than those reported in Point-MAE [26] and Point-BERT [55]. The results from our training-from-scratch baseline are presented in this section, as shown in Tab. 6, Tab. 7, and Tab. 8.
We follow the same evaluation protocol as Point-MAE [26]. While the results we obtained from the training-from-scratch baseline consistently surpass the previously reported scores, there remains a significant gap between the performance of the training-from-scratch baseline and the pre-trained methods. This underscores the effectiveness of pre-training in enhancing model performance. Furthermore, pre-trained methods demonstrate much faster convergence compared to the training-from-scratch baseline.
@@ Appendix B More Rigorous Evaluation
In prior works [55, 26], the validation set was identical to the test set. The model was evaluated on the test set after every epoch, with the best-performing result selected. Such practices can artificially inflate performance metrics and fail to accurately reflect the model’s ability to generalize to unseen data. While we followed this setup in the main text for fair comparisons, we also performed more rigorous evaluations using a dedicated validation set, which has no overlap with either the training set or the test sets. Specifically, we set aside 20% of the original test set as our validation set, leaving the remaining 80% as the new test set. We then report the test performance with checkpoints selected based on the validation accuracy.
@@
Scratch* [55] 92.95 84.19 86.94 80.92
Point-MAE-SN [26] 92.22 88.32 87.97 83.83
@@ Appendix C Linear Probing
In line with Point-BERT [55] and Point-MAE [26], we primarily report the performance of pre-trained methods in a transfer learning and few-shot learning setting in the main text, where the pre-trained model is fine-tuned with task-specific supervision. However, a common practice for benchmarking self-supervised learning methods is the linear probing, which trains a single linear layer while keeping the pretrained network backbone frozen. Below we provide details of our experimental setup and results.
@@ Appendix E Details on Generating Scene-Level Procedural 3D Programs
We follow MegaSynth’s  [17] generation pipeline. First, the pipeline generates a scene floor plan. Each 3D box represents a different shape, with distinct colors indicating various object types. This defines the spatial structure of the scene. Next, we creates 3D objects by combining primitive shapes such as cubes and spheres. These objects undergo geometry augmentations, including scaling, deformation, and boolean operations, to introduce variations. Since we do not need texture and lighting, we omit the rest of the procedurals. Instead we sample points from each shape and randomly sample points from the layout in order to represent walls, ceilings and floors. We will release code to further assist the reproducibility of our results.
@@
We closely follow each baseline’s open‐source configuration. Similar to prior work for 3D shape pretraining, we sample each point cloud to $p=1024$ points and divide it into $n=64$ patches, with each patch containing $k=32$ points via the KNN algorithm. The autoencoder consists of an encoder with 12 Transformer blocks and a decoder with 4 Transformer blocks, each block having a 384‐dimensional hidden size and 6 attention heads. During pretraining, we randomly sample 1024 points per shape and apply standard random scaling and translation. We train for 300 epochs using the AdamW optimizer [24] with a cosine decay schedule [23], an initial learning rate of 0.001, weight decay of 0.05, and a batch size of 128.
For scene-level SSL MSC, we use SparseUNet34 as the backbone with a hierarchical encoder-decoder structure. The encoder consists of depths $[2,3,4,6]$ and channels $[32,64,128,256]$, while the decoder follows depths $[2,2,2]$ with channels $[256,128,64,64]$. A kernel size of 3 is used in both encoding and decoding, with a pooling stride of $[2,2,2,2]$. The pre-training phase employs SGD [31] with a cosine decay scheduler [23], an initial learning rate of 0.1, a weight decay of $1e^{-4}$, and a momentum of 0.8. We use a batch size of $32$ and train on ScanNet for 600 epochs, with 6 warmup epochs. For fine-tuning, we use a similar SGD [31] with cosine decay setup [23], but with a learning rate of 0.05, weight decay of $1e^{-4}$, and momentum of 0.9. The batch size is 48, with 40 warmup epochs, and the model is trained for 600 epochs. We will release code to further assist the reproducibility of our results.
Figure 9: Masked Point Cloud Completion. This figure visualizes shape completion results with Point-MAE-SN and Point-MAE-Zero on the test split of ShapeNet and procedurally synthesized 3D shapes. Left: Ground truth 3D point clouds and masked inputs with a 60% mask ratio. Middle: Shape completion results using the centers of masked input patches as guidance, following the training setup of Point-MAE [26]. Right: Point cloud reconstructions without any guidance points. The $L_{2}$ Chamfer distance (lower is better) between the predicted 3D point clouds and the ground truth is displayed below each reconstruction.

### v3

@@ 4.5 Analysis
Fig. 4 displays samples from each configuration alongside quantitative comparisons of pretraining performance and downstream classification accuracy on PB-T50-RS, the most challenging variant of ScanObjectNN [uy2019revisiting]. As shape complexity increases, the pretraining task becomes more difficult, leading to higher reconstruction losses at the 300th training epoch. However, the downstream classification performance of Point-MAE-Zero improves. This underscores the importance of topological diversity in shapes for effective self-supervised point cloud representation learning.
@@
Figure 5: Impact of pretraining dataset size. We report the classification accuracy (%) on the PB-T50-RS subset of ScanObjectNN [uy2019revisiting] as a function of the pretraining dataset size.
@@
Dataset Size and Comparison with Objaverse. Fig. 5 presents a scaling analysis of Point-MAE-Zero on the PB-T50-RS benchmark using procedurally generated data. When matched for dataset size, Point-MAE-Zero and Point-MAE-SN achieve comparable downstream performance, underscoring the viability of synthetic data as a substitute for curated datasets like ShapeNet. Importantly, we observe that performance further improves with increasing dataset size, despite the lack of semantics in the synthetic data.
To contextualize our results, we compare against pretraining on randomly sampled subsets of Objaverse [deitke2023objaverse], matched in scale. While models pretrained on Objaverse exhibit strong performance at smaller scales (e.g., $10^{3}$ or $10^{4}$ shapes), their downstream performance plateaus—and eventually declines—as more shapes are added. We pretrain and finetune Point-MAE-Objaverse 5 times on PB-T50-RS and we report the best performance. We hypothesize that this trend is due to many randomly sampled shapes from Objaverse having simple geometry (e.g., avatars, boxes, cups), which makes the pretraining task too easy and limits the ability to learn robust 3D representations. In contrast, our procedurally generated data, enriched through augmentation and shape composition, consistently improves performance as the dataset scales, demonstrating its effectiveness and scalability for point cloud self-supervised learning.
Efficiency of Transfer Learning. Fig. 6 shows the learning curves for training from scratch, Point-MAE-SN, and Point-MAE-Zero on shape classification tasks in the transfer learning setting. Both Point-MAE-SN and Point-MAE-Zero demonstrate faster training convergence and higher accuracy compared to training from scratch, consistently across both ModelNet40 and ScanObjectNN benchmarks.
t-SNE Visualization. Fig. 7 visualizes the distribution of 3D shape representations from Point-MAE-SN and Point-MAE-Zero via t-SNE [van2008visualizing], before and after fine-tuning on specific downstream tasks. It also includes representations from a randomly initialized neural network as a reference.
First, compared to the representations from scratch, both Point-MAE-SN and Point-MAE-Zero demonstrate visually improved separation between different categories in the latent space. For example, this is evident in the red and light blue clusters on ModelNet40 and the blue and light blue clusters on ScanObjectNN. This highlights the effectiveness of masked auto-encoding for self-supervised 3D learning.
@@ 5 Discussion
This work has several limitations. For example, due to limited computational resources, we were unable to further scale up our experiments, such as by increasing the dataset size or conducting more detailed ablation studies on procedural 3D generation. Additionally, our findings may be influenced by potential biases in visualization tools (e.g., t-SNE) or benchmarks (e.g., data distribution and evaluation protocols). Furthermore, in 3D vision, the distinction between geometric structures and semantics remains an open question, as well-stated by Xie et al[xie2024lrm]. This work also does not provide any novel representation learning method. Nevertheless, we hope our findings will inspire further exploration into self-supervised 3D representation learning.
6 Acknowledgment
The authors acknowledge the Adobe Research Gift, the University of Virginia Research Computing and Data Analytics Center, Advanced Micro Devices AI and HPC Cluster Program, Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, and National Artificial Intelligence Research Resource (NAIRR) Pilot for computational resources, including the Anvil supercomputer (National Science Foundation award OAC 2005632) at Purdue University and the Delta and DeltaAI advanced computing resources (National Science Foundation award OAC 2005572).
[…]
@@ 6 Acknowledgment
We provide more implementation details, additional experimental results, and visualizations in this supplementary material. In Sec. A, we present our training-from-scratch baseline, which surpasses the results reported in Point-MAE [pang2022masked]. In Sec. B, we introduce a more rigorous evaluation protocol using a dedicated validation set instead of relying on the test set for validation. Moreover, in Sec. C, we provide linear probing results on ModelNet40 and three ScanObjectNN variants to compare the performance of Point-MAE-SN and Point-MAE-Zero. Additional visualizations are provided in Sec.D. Details on generating scene-level procedural 3D programs are in Sec.E. Lastly, implementation details for 3D object SSLs and 3D scene SSLs are presented in Sec. F.
@@ Appendix A Training-from-Scratch Baseline
For the training-from-scratch baseline, we report the results in the supplementary material for completeness. In the main text, we referenced the scores reported in the original Point-MAE paper for this baseline. However, our experiments suggest that training from scratch in our setup produced higher scores than those reported in Point-MAE [pang2022masked] and Point-BERT [yu2022point]. The results from our training-from-scratch baseline are presented in this section, as shown in Tab. 6, Tab. 7, and Tab. 8.
We follow the same evaluation protocol as Point-MAE [pang2022masked]. While the results we obtained from the training-from-scratch baseline consistently surpass the previously reported scores, there remains a significant gap between the performance of the training-from-scratch baseline and the pre-trained methods. This underscores the effectiveness of pre-training in enhancing model performance. Furthermore, pre-trained methods demonstrate much faster convergence compared to the training-from-scratch baseline.
@@ Appendix B More Rigorous Evaluation
In prior works [yu2022point, pang2022masked], the validation set was identical to the test set. The model was evaluated on the test set after every epoch, with the best-performing result selected. Such practices can artificially inflate performance metrics and fail to accurately reflect the model’s ability to generalize to unseen data. While we followed this setup in the main text for fair comparisons, we also performed more rigorous evaluations using a dedicated validation set, which has no overlap with either the training set or the test sets. Specifically, we set aside 20% of the original test set as our validation set, leaving the remaining 80% as the new test set. We then report the test performance with checkpoints selected based on the validation accuracy.
@@
Scratch* [yu2022point] 92.95 84.19 86.94 80.92
Point-MAE-SN [pang2022masked] 92.22 88.32 87.97 83.83
@@ Appendix C Linear Probing
In line with Point-BERT [yu2022point] and Point-MAE [pang2022masked], we primarily report the performance of pre-trained methods in a transfer learning and few-shot learning setting in the main text, where the pre-trained model is fine-tuned with task-specific supervision. However, a common practice for benchmarking self-supervised learning methods is the linear probing, which trains a single linear layer while keeping the pretrained network backbone frozen. Below we provide details of our experimental setup and results.
@@ Appendix E Details on Generating Scene-Level Procedural 3D Programs
We follow MegaSynth’s  [jiang2024megasynth] generation pipeline. First, the pipeline generates a scene floor plan. Each 3D box represents a different shape, with distinct colors indicating various object types. This defines the spatial structure of the scene. Next, we creates 3D objects by combining primitive shapes such as cubes and spheres. These objects undergo geometry augmentations, including scaling, deformation, and boolean operations, to introduce variations. Since we do not need texture and lighting, we omit the rest of the procedurals. Instead we sample points from each shape and randomly sample points from the layout in order to represent walls, ceilings and floors. We will release code to further assist the reproducibility of our results.
@@
We closely follow each baseline’s open‐source configuration. Similar to prior work for 3D shape pretraining, we sample each point cloud to $p=1024$ points and divide it into $n=64$ patches, with each patch containing $k=32$ points via the KNN algorithm. The autoencoder consists of an encoder with 12 Transformer blocks and a decoder with 4 Transformer blocks, each block having a 384‐dimensional hidden size and 6 attention heads. During pretraining, we randomly sample 1024 points per shape and apply standard random scaling and translation. We train for 300 epochs using the AdamW optimizer [loshchilov2016sgdr] with a cosine decay schedule [loshchilov2017decoupled], an initial learning rate of 0.001, weight decay of 0.05, and a batch size of 128.
For scene-level SSL MSC, we use SparseUNet34 as the backbone with a hierarchical encoder-decoder structure. The encoder consists of depths $[2,3,4,6]$ and channels $[32,64,128,256]$, while the decoder follows depths $[2,2,2]$ with channels $[256,128,64,64]$. A kernel size of 3 is used in both encoding and decoding, with a pooling stride of $[2,2,2,2]$. The pre-training phase employs SGD [ruder2016overview] with a cosine decay scheduler [loshchilov2017decoupled], an initial learning rate of 0.1, a weight decay of $1e^{-4}$, and a momentum of 0.8. We use a batch size of $32$ and train on ScanNet for 600 epochs, with 6 warmup epochs. For fine-tuning, we use a similar SGD [ruder2016overview] with cosine decay setup [loshchilov2017decoupled], but with a learning rate of 0.05, weight decay of $1e^{-4}$, and momentum of 0.9. The batch size is 48, with 40 warmup epochs, and the model is trained for 600 epochs. We will release code to further assist the reproducibility of our results.
Figure 9: Masked Point Cloud Completion. This figure visualizes shape completion results with Point-MAE-SN and Point-MAE-Zero on the test split of ShapeNet and procedurally synthesized 3D shapes. Left: Ground truth 3D point clouds and masked inputs with a 60% mask ratio. Middle: Shape completion results using the centers of masked input patches as guidance, following the training setup of Point-MAE [pang2022masked]. Right: Point cloud reconstructions without any guidance points. The $L_{2}$ Chamfer distance (lower is better) between the predicted 3D point clouds and the ground truth is displayed below each reconstruction.